Which pair has equally likely outcomes? Check the two choices below which have equal probabilities of success. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards).

rolling a sum of 4 on two fair six sided dice
drawing a black seven out of a standard 52 card deck given it’s not a face card or an Ace.
rolling a sum of 11 on two fair six sided dice
rolling a sum of 7 on two fair six sided dice
drawing a three out of a standard 52 card deck given it’s not a face card or an Ace

the fractions I come up with no matches
1/13 drawing 3 out of deck of cards
2/52 drawing a black 7 out of deck of cards
6/36 = 1/6 rolling 7
2/36 rolling 11

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

rolling a sum of 4 on two fair six sided dice:::3/36 = 1/12
------
drawing a black seven out of a standard 52 card deck given it’s not a face card or an Ace. ::: 2/36 = 1/18
------
rolling a sum of 11 on two fair six sided dice:::2/36 = 1/18
-------
rolling a sum of 7 on two fair six sided dice:::1/6
--------
drawing a three out of a standard 52 card deck
given it’s not a face card or an Ace:::4/36 = 1/9
-----------
Cheers,
Stan H.
=============
1/13 drawing 3 out of deck of cards
2/52 drawing a black 7 out of deck of cards
6/36 = 1/6 rolling 7
2/36 rolling 11
=============

Answer by ewatrrr(24785)


Hi,
Must take into account the given when considering the card questions.
Otherwise You are on the right track!!
GIVEN it's not a face card or an ace...
drawing a black seven out of a standard 52 card deck
52 - 4aces - 12face cards = 36 possibilities P = 2/36 = 1/18 *
drawing a 3 out of a standard 52 card deck 4/36 = 1/9
rolling a sum of 4 on two fair six sided dice. P = 3/36 = 1/12
rolling a sum of 11 on two fair six sided dice. P = 2/36 = 1/18 *
rolling a sum of 7 on two fair six sided dice. P = 6/36 = 1/6
those starred have equal probabilities of success
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12